OpenStudy (samsan9):

find the cente, focus, directrix, and length of the focal chord for y^2+6y-2x+13=0.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ y^2+6y-2x+13=0 }\) \(\normalsize\color{blue}{ y^2+6y-2x+9+4=0 }\) \(\normalsize\color{blue}{ y^2+6y+9-2x+4=0 }\) \(\normalsize\color{blue}{ (y-9)^2-2x+4=0 }\) \(\normalsize\color{blue}{ (y-9)^2+4=x }\) \(\normalsize\color{blue}{ x=(y-9)^2+4 }\) same thing as \(\normalsize\color{blue}{ x=1(y-9)^2+4 }\) horizontal parabola.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x=a(y-h)^2+k }\) (h,k) is the center